OpenStudy (ashleyisakitty):

If the scale factor of two similar solids is 4:9, what is the ratio of their corresponding areas? What is the ratio of their corresponding volumes? Please help, thank you.

OpenStudy (ashleyisakitty):

@Compassionate would you know how to do this problem?

OpenStudy (compassionate):

So, for the 4:9 we have 4/9 We just square it 16:81 16/81 = 0.2 I have to hurry in the shower, be back in a minute. Need to do it before my relative goes to bed. If you're still here :)

OpenStudy (ashleyisakitty):

is that volume or area?

OpenStudy (ashleyisakitty):

Okay ill be here waiting :)

OpenStudy (compassionate):

I'm back!

OpenStudy (ashleyisakitty):

Great, so were your calculations the volume ratio?

OpenStudy (compassionate):

Alright, one second, I was calculating the ratio. Ha ha. Okay, give me a minute to re-do it.

OpenStudy (ashleyisakitty):

16:81 is correct but did you do it for volume or area?

OpenStudy (compassionate):

That's the ratio of the solids

OpenStudy (ashleyisakitty):

Oh.

OpenStudy (compassionate):

Take the scale factor and square it to get the ratio.

OpenStudy (ashleyisakitty):

well the area or volume will be 16:81

OpenStudy (compassionate):

Right, Nicole. That's true. For areas, just square the ratio. (4:9)^2 = (16:81) For volumes, just cube. (4:9)^3 = (64:729)

OpenStudy (ashleyisakitty):

Right. Thats what I thought. Thank you.

OpenStudy (compassionate):

Lets jump over to your other question.